(2x-2)^2=10
How do you solve finding the square root?

Answer by tommyt3rd(5050)

(2x-2)^2=10
(2x-2)=sqrt(10), or (2x-2)=-sqrt(10)
x=1+(1/2)sqrt(10) or x=1-(1/2)sqrt(10)